OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. 2x + 4y - 3z = -7 3x + y + 4z = -12 x + 3y + 4z = 4

OpenStudy (anonymous):

a) (-6, -2, -1) b) (6, -2, -1) c) (-6, 2, -1) d) (-6, 2, 1)

OpenStudy (jdoe0001):

why not just solve it? :)

OpenStudy (jdoe0001):

pick 2 above, cancel out "z" pick another 2 from above including the non-included one before, cancel "z" use the 2 new resultant equations solve for either "x" or "y" then extrapolate to the other equations

OpenStudy (anonymous):

extrapolate? i dont get how to even set up the question... i dont think i can solve it.

OpenStudy (jdoe0001):

well, pick the 1st two, 2x + 4y - 3z = -7 3x + y + 4z = -12 and cancel "z"

OpenStudy (jdoe0001):

then pick another 2, including the one you didn't use previously, say 1st and 3rd equations 2x + 4y - 3z = -7 x + 3y + 4z = 4 and cancel "z" again

OpenStudy (jdoe0001):

so, you'd get a RESULTANT equation from the 1st cancellation, and another RESULTANT equation from the 2nd cancellation those 2 RESULTANT equations, will just have "x" and "y", then you'd just solve for either, by eliminating the other

OpenStudy (anonymous):

okay ill give it a try

OpenStudy (jdoe0001):

2x + 4y - 3z = -7 times 4 3x + y + 4z = -12 times 3 8x + 16y - 12z = -28 9x + 3y + 12z = -36 ------------------- 17x+19y+0 =- 64

OpenStudy (jdoe0001):

2x + 4y - 3z = -7 times 4 x + 3y + 4z = 4 times 3 8x + 16y - 12z = -28 3x + 9y + 12z = 12 -------------------- 11x+25y+ 0 = -16

OpenStudy (anonymous):

just to let you know...you wouldn't have to cancel out z....you can cancel out any letters x or y also.

OpenStudy (jdoe0001):

and just eliminate from there on those 2, "without a z" equations, and solve for either, x or y

OpenStudy (anonymous):

just remember, whatever you cancel out on the first two equations, you have to cancel the same letter out on the last two equations.

OpenStudy (jdoe0001):

indeed

OpenStudy (anonymous):

okay i tried them all i think i got answer d) (-6, 2, 1)

OpenStudy (anonymous):

2x + 4y - 3z = - 7 3x + y + 4z = - 12 -->(-4)3x + y + 4z = -12 ----------------- 2x + 4y - 3z = - 7 -12x - 4y - 16z = 48 (result of multiplying by -4) ----------------- -10x + 0y - 19z = 41 -10x - 19z = 41 3x + y + 4z = - 12 -->(-3)3x + y + 4z = - 12 x + 3y + 4z = 4 ------------------ -9x - 3y - 12z = 36 (result of multiplying by -3) x + 3y + 4z = 4 -----------------add -8x + 0 - 8z = 40 -8x -8z = 40 -10x - 19z = 41 -->(4)-10x - 19z = 41 -8x - 8z = 40 -->(-5) -8x - 8z = 40 ---------------- -40x - 76z = 164 (result of multiplying by 4) 40x + 40z = - 200 (result of multiplying by -5) -----------------add 0 - 36z = - 36 -36z = - 36 z = -36/-36 = 1 now sub 1 in for z -8x - 8z = 40 -8x - 8(1) = 40 -8x -8 = 40 -8x = 40 + 8 -8x = 48 x = - 48/8 x = - 6 now sub in known values in any of the original equations 2x + 4y - 3z = - 7 2(-6) + 4y - 3(1) = - 7 -12 + 4y - 3 = - 7 4y = - 7 + 12 + 3 4y = 8 y = 8/4 y = 2 check.... 3x + y + 4z = - 12 3(-6) + 2 + 4(1) = - 12 -18 + 2 + 4 = - 12 -18 + 6 = - 12 -12 = - 12 (correct) you got it......answer is (-6,2,1)